a 2000 lb block lies on a smooth plaane at 50 degrees with the horizontal. Find the horizontal force needed to hold the block on the inclined plane.


1
Expert's answer
2021-03-25T19:28:44-0400

Let's apply the Newton's Second Law of Motion:


"F_h-mgsin\theta=0,""F_h=mgsin\theta,""F_h=2000\ lb\cdot\dfrac{0.453592\ kg}{1\ lb}\cdot9.8\ \dfrac{m}{s^2}\cdot sin50^{\circ}=6810\ N."